Your factory manufactures tricycles. You need to purchase front wheels from the supplier. The expected annual demand (D) is 1300 units; the setup or ordering cost (S) is $8 per order, and the holding cost per unit per year (H) is $0.50.  

  1. Determine optimal number of wheels to order
  2. Determine the reorder point (ROP) (without safety stock) assuming that the lead time (L) is 4 days and your factory works 250 days a year.
  3. Based on your answer in (b), determine the reorder point (ROP) (with safety stock) assuming that the standard deviation of demand during the lead time (σdLT) is 10 and acceptable service level is 95%. (Z for 95% service level in 1.65)
Answer #1

1. DEMAND = 1300

ORDERING COST = 8

HOLDING COST = 0.5

EOQ = SQRT(2 * DEMAND * ORDERING COST / HOLDING COST) = SQRT(2 * 1300 * 8 / 0.5) = 204

2. DAILY DEMAND = DEMAND / NO. OF WORKING DAYS = 1300 / 250 = 5.2

REORDER POINT = DAILY DEMAND * LEAD TIME = 5.2 * 4 = 20.8

3. SERVICE LEVEL = 95, WHICH GIVES A Z VALUE OF 1.645

REORDER POINT = DEMAND DURING LEAD TIME + SAFETY STOCK

SAFETY STOCK = Z * STANDARD DEVIATION OF DEMAND DURING LEAD TIME

SAFETY STOCK = 1.645 * 10 = 16.5

REORDER POINT = 20.8 + 16.5 = 37